# supplementary[supplementary.pdf]

### v1

@@ 6 Experiments
We evaluate our proposed method through POMDP simulations conducted in three distinct settings. Firstly, we investigate the exact computation of the original POMDP and the proposed bounds to validate the findings of the theoretical analysis presented in Section 4. Secondly, we assess a sparse sampling POMDP solver using our estimated bounds, as outlined in Section 5, addressing larger POMDP problems. Finally, we apply the sparse sampling method to a beacon navigation problem to showcase the potential of our approach in practical robotics applications. The detailed experiment settings appear in Section LABEL:sec:experiment-settings of the Supplementary document [17].

### v2

@@ 6 Experiments
We evaluate our proposed method through POMDP simulations conducted in three distinct settings. Firstly, we investigate the exact computation of the original POMDP and the proposed bounds to validate the findings of the theoretical analysis presented in Section 4. Secondly, we assess a sparse sampling POMDP solver using our estimated bounds, as outlined in Section 5, addressing larger POMDP problems. Finally, we apply the sparse sampling method to a beacon navigation problem to showcase the potential of our approach in practical robotics applications. The detailed experiment settings appear in Section F of the Supplementary document [17].